# Learning-Augmented Priority Queues

### v1

@@ Abstract
Priority queues are one of the most fundamental and widely used data structures in computer science. Their primary objective is to efficiently support the insertion of new elements with assigned priorities and the extraction of the highest priority element. In this study, we investigate the design of priority queues within the learning-augmented framework, where algorithms use potentially inaccurate predictions to enhance their worst-case performance. We examine three prediction models spanning different use cases, and show how the predictions can be leveraged to enhance the performance of priority queue operations. Moreover, we demonstrate the optimality of our solution and discuss some possible applications.
@@ 1.1 Problem definition
This prediction model was introduced in [Bai and Coester, 2023] for sorting, but it has a broader significance in comparison-based problems, such as search [Borgstrom and Kosaraju, 1993; Nowak, 2009; Tschopp et al., 2011], ranking [Wauthier et al., 2013; Shah and Wainwright, 2018; Heckel et al., 2018], and the design of comparison-based ML algorithms [Haghiri et al., 2017, 2018; Ghoshdastidar et al., 2019; Perrot et al., 2020; Meister and Nietert, 2021]. Particularly, it has great theoretical importance for priority queues, which have been extensively studied in the comparison-based framework [Gonnet and Munro, 1986; Brodal and Okasaki, 1996; Edelkamp and Wegener, 2000]. Comparison-based models are often used, for example, when the preferences are determined by human subjects. Assigning numerical scores in these cases is inexact and prone to errors, while pairwise comparisons are absolute and more robust [David, 1963]. Within such a setting, dirty comparisons can be obtained by a binary classifier, and used to minimize human inference yielding clean comparisons, which are time-consuming and might incur additional costs.
@@
The last setting assumes that the priority queue is used in a process where a finite number $N$ of distinct keys will be inserted, i.e., the priority queue is not used indefinitely, but $N$ is unknown to the algorithm. Upon the insertion of any new key $u_{i}$, the algorithm receives a prediction $\widehat{R}(u_{i})$ of the rank of $u_{i}$ within all the $N$ keys $\{u_{j}\}_{j\in[N]}$. Denoting by $R(u_{i})=r(u_{i},\{u_{j}\}_{j\in[N]})$ the true rank, the prediction error of $u_{i}$ is
@@
Note that $N$ counts the distinct keys that are added either with $\operatorname{Insert}$ or $\operatorname{DecreaseKey}$ operations. An arbitrarily large number of $\operatorname{DecreaseKey}$ operations thus can make $N$ arbitrarily large although the total number of insertions is reduced. However, with a lazy implementation of $\operatorname{DecreaseKey}$, we can assume without loss of generality that $N$ is at most quadratic in the total number of insertions. Indeed, it is possible to omit executing the $\operatorname{DecreaseKey}$ operations when queried, and only store the new elements’ keys to update. Then, at the first $\operatorname{ExtractMin}$ operation, all the element’s keys are updated by executing for each one only the last queried $\operatorname{DecreaseKey}$ operation involving it. Denoting by $k$ the total number of insertions, there are at most $k$ $\operatorname{ExtractMin}$ operations, and for each of them, there are at most $k$ $\operatorname{DecreaseKey}$ operations executed. The total number of effectively executed $\operatorname{DecreaseKey}$ operations is therefore $O(k^{2})$.
@@ 1.2 Our results
We first investigate augmenting binary heaps with predictions. A binary heap $\mathcal{Q}$ with $n$ elements necessitates $O(\log n)$ time but only $O(\log\log n)$ comparisons for insertion [Gonnet and Munro, 1986]. To leverage dirty comparisons, we first design a randomized binary search algorithm to find the position of an element $u$ in a sorted list $L$. We prove that it terminates using $O(\#L)$ dirty comparisons and $O(\log\eta(u,L))$ clean comparisons in expectation. Subsequently, we use this result to establish an insertion algorithm in binary heaps using $O(\log\log n)$ dirty comparisons and reducing the number of clean comparisons to $O(\log\log\eta(u,\mathcal{Q}))$ in expectation. However, $\operatorname{ExtractMin}$ still mandates $O(\log n)$ clean comparisons. In the two other prediction models, binary heaps and other heap implementations of priority queues appear unsuitable, as the positions of the keys are not determined solely by their ranks.
Consequently, in Section 3, we shift to using skip lists. We devise randomized insertion algorithms requiring, in expectation, only $O(\log\vec{\eta}(u,\mathcal{Q}))$ time in the pointer prediction model, $O(\log\eta(u,\mathcal{Q}))$ clean comparisons in the dirty comparison model, and $O(\log\log N+\log\max_{i\in[N]}\eta^{\Delta}(u_{i}))$ time in the rank prediction model, where we use in the latter an auxiliary van Emde Boas (vEB) tree [van Emde Boas et al., 1976]. Note that, with perfect rank predictions, the insertion runtime decreases to $O(\log\log N)$. Across the three prediction models, $\operatorname{FindMin}$ and $\operatorname{ExtractMin}$ only necessitate $O(1)$ time, and the complexity of $\operatorname{DecreaseKey}$ aligns with that of insertion. Finally, we prove in Theorem 4.1 the optimality of our data structure.
Our learning-augmented data structure enables additional operations beyond those of priority queues, such as the maximum priority queue operations $\operatorname{FindMax}$, $\operatorname{ExtractMax}$, and $\operatorname{IncreaseKey}$ with analogous complexities, removing an arbitrary key $u$ from the priority queue in expected $O(1)$ time, and finding the predecessor or successor of any key $u$ belonging to the priority queue.
@@ 1.3 Related work
The dirty and clean comparison model is also related to the prediction querying model, gaining a growing interest in the study of learning-augmented algorithms, where the decision-maker decides when to query predictions and for which items [Im et al., 2022; Benomar and Perchet, 2024a; Sadek and Elias, 2024]. In particular, having free-weak and costly-strong oracles has been studied in [Silwal et al., 2023] for correlation clustering, in [Bateni et al., 2023] for clustering and computing minimum spanning trees in a metric space, and in [Eberle et al., 2024] for matroid optimization. Another related setting involves the algorithm observing partial information online, which it can then use to decide whether to query a costly hint about the current item. This has been explored in contexts such as online linear optimization [Bhaskara et al., 2021] and the multicolor secretary problem [Benomar et al., 2023].
@@ 3 Skip lists
A priority queue can be implemented naively by maintaining a dynamic sorted linked list of keys. This guarantees constant time for $\operatorname{ExtractMin}$, but $O(n)$ time for insertion. Skip lists offer a solution to this inefficiency, by maintaining multiple levels of linked lists, with higher levels containing fewer elements and acting as shortcuts to lower levels, facilitating faster search and insertion in expected $O(\log n)$ time. In all subsequent discussions concerning linked lists or skip lists, it is assumed that they are doubly linked, having both predecessor and successor pointers between elements.
@@
The keys of the priority queue can be considered pairwise distinct by grouping elements with the same key together in a collection, for example, a HashSet. This collection can be accessible in $O(1)$ time via the key using a HashMap. When a new item $x$ with priority $u$ is to be inserted, if $u$ is already in the priority queue, then $x$ is added to the corresponding collection in $O(1)$ time. With such implementation, when an $\operatorname{ExtractMin}$ operation is called, if multiple elements correspond to the minimum key, then the algorithm can for example retrieve an arbitrary one of them, of the first inserted one, depending on the use case.
@@
Before presenting the insertion algorithms in the three prediction models, we prove an upper bound on the expected maximum of i.i.d. geometric random variables with parameter $p$. This Lemma will be useful in the analysis of our algorithms, as the heights of elements in the skip list are i.i.d. geometric random variables.
Lemma 3.1.
If $X_{1},\ldots,X_{m}$ are i.i.d. random variables following a geometric random distribution with parameter $p$, then
$\mathbb{E}[\max_{i\in[m]}X_{i}]\leq\log_{\frac{1}{1-p}}(m)+\tfrac{1}{\ln 1/q}+1=O(\log(m+1))\;.$
Proof.
Let $q=1-p$ and $Z=\max_{i\in[m]}X_{i}$. It holds for all $k\geq 1$ that $\Pr(Z_{1}\leq k)=1-q^{k}$, hence $\Pr(Z\leq k)=(1-q^{k})^{m}$, and it follows that
$\mathbb{E}[Z]$ $=\sum_{k=1}^{\infty}\Pr(Z\geq k)=1+\sum_{k=2}^{\infty}(1-\Pr(Z\leq k-1))$
$=1+\sum_{k=1}^{\infty}(1-(1-q^{k})^{m})\leq 1+\int_{0}^{\infty}(1-(1-q^{x})^{m})\mathrm{d}x\;,$
and setting $y=1-q^{x}$, i.e. $x=\frac{\ln(1-y)}{\ln q}$, we obtain
$\mathbb{E}[Z]$ $\leq 1+\int_{0}^{1}(1-y^{m})\left(\frac{-1}{(\ln q)(1-y)}\right)$
$=1+\frac{1}{\ln 1/q}\int_{0}^{\infty}\frac{1-y^{m}}{1-y}\mathrm{d}y$
$=1+\frac{1}{\ln 1/q}\sum_{k=0}^{m-1}\int_{0}^{1}y^{k}\mathrm{d}y=1+\frac{1}{\ln 1/q}\sum_{k=1}^{m}\frac{1}{k}$
$\leq 1+\frac{1}{\ln 1/q}\left(1+\ln m\right)=\log_{1/q}m+\tfrac{1}{\ln 1/q}+1\;.$
@@ 3.1 Pointer prediction
Assuming that $v_{j}<u$, we describe an algorithm for finding the predecessor of $u$ starting from the position of the key $v_{j}$ in the priority queue. If $v_{j}>u$, the search direction is reversed.
@@ 8      $\ell\leftarrow\ell-1$;
The number of comparisons made in this phase is therefore at most the number of comparisons needed to reach level $10pt^{*}(v_{j},u)+$ 1 starting from $v_{j}$ using the Bottom-Up search. We consider the hypothetical setting where the skip list is infinite to the right, the expected number of comparisons to reach level $10pt^{*}(v_{j},u)+$ 1 in this case is an upper bound on the expected number of comparisons needed in the Bottom-Up phase of Algorithm 2, as the algorithm also terminates if the end of the skip-list is reached. Let $T(\ell)$ be the expected number of comparisons made in the bottom-up search to reach level $\ell$ in an infinite skip-list. After each comparison made in the bottom-up search, it is possible to go at least one level up with probability $p$, while the algorithm can only move horizontally to the right with probability $1-p$. This induces the inequality
@@ 3.3.1 Van Emde Boas (vEB) trees
A vEB tree over an interval $\{i+1,\ldots,i+m\}$ has a root with $\sqrt{m}$ children, each being the root of a smaller vEB tree over a sub-interval $\{i+k\sqrt{m}+1,\ldots,i+(k+1)\sqrt{m}\}$ for some $k\in\{0,\ldots,\sqrt{m}-1\}$. The tree leaves are either empty or contain elements with the corresponding key, stored together in a collection, and internal nodes carry binary information indicating if the subtree they root contains, or not, at least one element. Denoting by $H(m)$ the height of a vEB tree of size $m$, it holds that $H(m)=1+H(\sqrt{m})$, which yields that $H(m)=O(\log\log m)$, enabling efficient implementation of the operations listed below in $O(\log\log m)$ time
@@ 4 Lower bounds
The following theorem gives lower bounds on the complexities of $\operatorname{ExtractMin}$ and $\operatorname{Insert}$.
@@
For each of the three prediction models, we establish the following lower bounds
@@
    Dirty comparisons: no data structure $\mathcal{Q}$ can support $\operatorname{ExtractMin}$ with $O(1)$ clean comparisons and $\operatorname{Insert}(u)$ with $o(\log\eta(u,\mathcal{Q}))$ clean comparisons.
@@
    Pointer predictions: no data structure $\mathcal{Q}$ can support $\operatorname{ExtractMin}$ with $O(1)$ comparisons and $\operatorname{Insert}(u)$ with $o(\log\vec{\eta}(u,\mathcal{Q}))$ comparisons.
@@
    Rank predictions: no data structure $\mathcal{Q}$ can support $\operatorname{ExtractMin}$ with $O(1)$ comparisons and $\operatorname{Insert}(u_{i})$ with $o(\log\max_{i\in[N]}\eta^{\Delta}(u_{i}))$ comparisons for all $i\in[N]$.
@@ 5.1 Sorting algorithm
for all $i\in\{2,\ldots,n\}$. By Theorem 3.2, the runtime for inserting $a_{\pi(i)}$ using this pointer prediction is $O(\log\vec{\eta}(a_{\pi(i)},\mathcal{Q}^{i-1}))$ The expected total time for inserting all the elements into the priority queue is therefore at most proportional to

### v2

@@ Abstract
Priority queues are one of the most fundamental and widely used data structures in computer science. Their primary objective is to efficiently support the insertion of new elements with assigned priorities and the extraction of the highest priority element. In this study, we investigate the design of priority queues within the learning-augmented framework, where algorithms use potentially inaccurate predictions to enhance their worst-case performance. We examine three prediction models spanning different use cases, and we show how the predictions can be leveraged to enhance the performance of priority queue operations. Moreover, we demonstrate the optimality of our solution and discuss some possible applications.
@@ 1.1 Problem definition
This prediction model was introduced in [Bai and Coester, 2023] for sorting, but it has a broader significance in comparison-based problems, such as search [Borgstrom and Kosaraju, 1993; Nowak, 2009; Tschopp et al., 2011], ranking [Wauthier et al., 2013; Shah and Wainwright, 2018; Heckel et al., 2018; El Ferchichi et al., ], and the design of comparison-based ML algorithms [Haghiri et al., 2017, 2018; Ghoshdastidar et al., 2019; Perrot et al., 2020; Meister and Nietert, 2021]. Particularly, it has great theoretical importance for priority queues, which have been extensively studied in the comparison-based framework [Gonnet and Munro, 1986; Brodal and Okasaki, 1996; Edelkamp and Wegener, 2000]. Comparison-based models are often used, for example, when the preferences are determined by human subjects. Assigning numerical scores in these cases is inexact and prone to errors, while pairwise comparisons are absolute and more robust [David, 1963]. Within such a setting, dirty comparisons can be obtained by a binary classifier, and used to minimize human inference yielding clean comparisons, which are time-consuming and might incur additional costs.
@@
The last setting assumes that the priority queue is used in a process where a finite number $N$ of distinct keys will be inserted, i.e., the priority queue is not used indefinitely, but $N$ is unknown to the algorithm. Upon the insertion of any new key $u_{i}$, the algorithm receives a prediction $\widehat{R}(u_{i})$ of the rank of $u_{i}$ among all the $N$ keys $\{u_{j}\}_{j\in[N]}$. Denoting by $R(u_{i})=r(u_{i},\{u_{j}\}_{j\in[N]})$ the true rank of $u_{i}$, the prediction error of $u_{i}$ is
@@
Note that $N$ counts the distinct keys that are added either with $\operatorname{Insert}$ or $\operatorname{DecreaseKey}$ operations. An arbitrarily large number of $\operatorname{DecreaseKey}$ operations thus can make $N$ arbitrarily large although the total number of insertions is reduced. However, with a lazy implementation of $\operatorname{DecreaseKey}$, we can assume without loss of generality that $N$ is at most quadratic in the total number of insertions. Indeed, it is possible to omit executing the $\operatorname{DecreaseKey}$ operations when queried, and only store the new elements’ keys to update. Then, at the first $\operatorname{ExtractMin}$ operation, all the element’s keys are updated by executing for each one only the last queried $\operatorname{DecreaseKey}$ operation involving it. Denoting by $k$ the total number of insertions, there are at most $k$ $\operatorname{ExtractMin}$ operations, and for each of them, there are at most $k$ $\operatorname{DecreaseKey}$ operations executed. The total number of effectively executed $\operatorname{DecreaseKey}$ operations is therefore $O(k^{2})$. In particular, this implies that a complexity of $O(\log N)$ is also logarithmic in the total number of insertions.
@@ 1.2 Our results
We first investigate augmenting binary heaps with predictions. A binary heap $\mathcal{Q}$ with $n$ elements necessitates $O(\log n)$ time but only $O(\log\log n)$ comparisons for insertion [Gonnet and Munro, 1986]. To leverage dirty comparisons, we first design a randomized binary search algorithm to find the position of an element $u$ in a sorted list $L$. We prove that it terminates using $O(\log(\#L))$ dirty comparisons and $O(\log\eta(u,L))$ clean comparisons in expectation. Subsequently, we use this result to establish an insertion algorithm in binary heaps using $O(\log\log n)$ dirty comparisons and reducing the number of clean comparisons to $O(\log\log\eta(u,\mathcal{Q}))$ in expectation. However, $\operatorname{ExtractMin}$ still mandates $O(\log n)$ clean comparisons. In the two other prediction models, binary heaps and other heap implementations of priority queues appear unsuitable, as the positions of the keys are not determined solely by their ranks.
Consequently, in Section 3, we shift to using skip lists. We devise randomized insertion algorithms requiring, in expectation, only $O(\log\vec{\eta}(u,\mathcal{Q}))$ time and comparisons in the pointer prediction model, $O(\log n)$ time and $O(\log\eta(u,\mathcal{Q}))$ clean comparisons in the dirty comparison model, and $O(\log\log N+\log\max_{i\in[N]}\eta^{\Delta}(u_{i}))$ amortized time and $O(\log\max_{i\in[N]}\eta^{\Delta}(u_{i}))$ comparisons in the rank prediction model, where we use in the latter an auxiliary van Emde Boas (vEB) tree [van Emde Boas et al., 1976]. Across the three prediction models, $\operatorname{FindMin}$ and $\operatorname{ExtractMin}$ only necessitate $O(1)$ time, and the complexity of $\operatorname{DecreaseKey}$ aligns with that of insertion. Finally, we prove in Theorem 4.1 the optimality of our data structure. Table 1 summarizes the complexities of our learning-augmented priority queue (LAPQ) in the three prediction models compared to standard priority queue implementations.
Priority queues $\operatorname{FindMin}$ $\operatorname{ExtractMin}$ $\operatorname{Insert}$ $\operatorname{DecreaseKey}$
Binary Heap $O(1)$ $O(\log n)$ $O(\log\log n)$ $O(\log n)$
Fibonacci Heap (amortized) $O(1)$ $O(\log n)$ $O(1)$
Skip List (average) $O(1)$ $O(1)$ $O(\log n)$
LAPQ with dirty comparisons (average) $O(1)$ $O(1)$ $O(\log\eta(u,\mathcal{Q}))$
LAPQ with pointer predictions (average) $O(1)$ $O(1)$ $O(\log\vec{\eta}(u,\mathcal{Q}))$
LAPQ with rank predictions (average) $O(1)$ $O(1)$ $O(\log\max_{i\in[n]}\eta^{\Delta}(u_{i}))$
Table 1: Number of comparisons per operation used by different priority queues.
Our learning-augmented data structure enables additional operations beyond those of priority queues, such as the maximum priority queue operations $\operatorname{FindMax}$, $\operatorname{ExtractMax}$, and $\operatorname{IncreaseKey}$ with analogous complexities, and removing an arbitrary key $u$ from the priority queue, finding its predecessor or successor in expected $O(1)$ time.
@@ 1.3 Related work
The dirty and clean comparison model is also related to the prediction querying model, gaining a growing interest in the study of learning-augmented algorithms, where the decision-maker decides when to query predictions and for which items [Im et al., 2022; Benomar and Perchet, 2024a; Sadek and Elias, 2024]. In particular, having free-weak and costly-strong oracles has been studied in [Silwal et al., 2023] for correlation clustering, in [Bateni et al., 2023] for clustering and computing minimum spanning trees in a metric space, and in [Eberle et al., 2024] for matroid optimization. Another related setting involves the algorithm observing partial information online, which it can then use to decide whether to query a costly hint about the current item. This has been explored in contexts such as online linear optimization [Bhaskara et al., 2021] and the multicolor secretary problem [Benomar et al., 2024].
@@ 3 Skip lists
A priority queue can be implemented naively by maintaining a dynamic sorted linked list of keys. This guarantees constant time for $\operatorname{ExtractMin}$, but $O(n)$ time for insertion. Skip lists (see Figure 1) offer a solution to this inefficiency, by maintaining multiple levels of linked lists, with higher levels containing fewer elements and acting as shortcuts to lower levels, facilitating faster search and insertion in expected $O(\log n)$ time. In all subsequent discussions concerning linked lists or skip lists, it is assumed that they are doubly linked, having both predecessor and successor pointers between elements.
@@
The keys of the priority queue can be considered pairwise distinct by grouping elements with the same key together in a collection, for example, a HashSet. This collection can be accessible in $O(1)$ time via the key using a HashMap. When a new item $x$ with priority $u$ is to be inserted, if $u$ is already in the priority queue, then $x$ is added to the corresponding collection in $O(1)$ time. With such implementation, when an $\operatorname{ExtractMin}$ operation is called, if multiple elements correspond to the minimum key, then the algorithm can, for example, retrieve an arbitrary one of them, or the first inserted one, depending on the use case. In the following, we present separately for each model an insertion algorithm leveraging the predictions.
@@
Before presenting the insertion algorithms in the three prediction models, we present an upper bound from [Eisenberg, 2008] on the expected maximum of i.i.d. geometric random variables with parameter $p$. This Lemma will be useful in the analysis of our algorithms, as the heights of elements in the skip list are i.i.d. geometric random variables.
Lemma 3.1 ([Eisenberg, 2008]).
If $X_{1},\ldots,X_{m}$ are i.i.d. random variables following a geometric random distribution with parameter $p$, then, denoting by $q=1-p$, it holds that
$\mathbb{E}[\max_{i\in[m]}X_{i}]\leq 1+\frac{1}{\log(1/q)}\sum_{k=1}^{m}\frac{1}{k}=O(\log_{1/q}m)\;.$
@@ 3.1 Pointer prediction
Given a pointer prediction $v_{j}=\widehat{\operatorname{Pred}}(u,\mathcal{Q})$, we describe below an algorithm for finding the true predecessor of $u$ starting from the position of the key $v_{j}$, then inserting $u$. We assume in the algorithm that $v_{j}\leq u$. If $v_{j}>u$, then the algorithm can be easily adapted by reversing the search direction.
@@ 8      $\ell\leftarrow\ell-1$;
The number of comparisons made in this phase is therefore at most the number of comparisons needed to reach level $10pt^{*}(v_{j},u)+$ 1 starting from $v_{j}$ using the Bottom-Up search. We consider the hypothetical setting where the skip list is infinite to the right, the expected number of comparisons to reach level $10pt^{*}(v_{j},u)+$ 1, in this case, is an upper bound on the expected number of comparisons needed in the Bottom-Up phase of Algorithm 2, as the algorithm also terminates if the end of the skip-list is reached. Let $T(\ell)$ be the expected number of comparisons made in the bottom-up search to reach level $\ell$ in an infinite skip list. After each comparison made in the bottom-up search, it is possible to go at least one level up with probability $p$, while the algorithm can only move horizontally to the right with probability $1-p$. This induces the inequality
@@ 3.3.1 Van Emde Boas (vEB) trees
A vEB tree over an interval $\{i+1,\ldots,i+m\}$ has a root with $\sqrt{m}$ children, each being the root of a smaller vEB tree over a sub-interval $\{i+k\sqrt{m}+1,\ldots,i+(k+1)\sqrt{m}\}$ for some $k\in\{0,\ldots,\sqrt{m}-1\}$. The tree leaves are either empty or contain elements with the corresponding key, stored together in a collection, and internal nodes carry binary information indicating whether or not the subtree they root contains at least one element. Denoting by $H(m)$ the height of a vEB tree of size $m$, it holds that $H(m)=1+H(\sqrt{m})$, which yields that $H(m)=O(\log\log m)$, enabling efficient implementation of the operations listed below in $O(\log\log m)$ time:
@@ 4 Lower bounds
As explained earlier, $\operatorname{ExtractMin}$ requires only $O(1)$ expected time in skip lists. Furthermore, we presented insertion algorithms for the three prediction models and provided upper bounds on their complexities. The following theorem establishes lower bounds on the complexities of $\operatorname{ExtractMin}$ and $\operatorname{Insert}$ for any priority queue augmented with any of the three prediction types.
@@
For each of the three prediction models, the following lower bounds hold.
@@
    Dirty comparisons: no data structure $\mathcal{Q}$ can support $\operatorname{ExtractMin}$ with $O(1)$ clean comparisons and $\operatorname{Insert}(u)$ with $o(\log\eta(u,\mathcal{Q}))$ clean comparisons in expectation.
@@
    Pointer predictions: no data structure $\mathcal{Q}$ can support $\operatorname{ExtractMin}$ with $O(1)$ comparisons and $\operatorname{Insert}(u)$ with $o(\log\vec{\eta}(u,\mathcal{Q}))$ comparisons in expectation.
@@
    Rank predictions: no data structure $\mathcal{Q}$ can support $\operatorname{ExtractMin}$ with $O(1)$ comparisons and $\operatorname{Insert}(u_{i})$ with $o(\log\max_{i\in[N]}\eta^{\Delta}(u_{i}))$ comparisons in expectation, for all $i\in[N]$.
@@ 5.1 Sorting algorithm
for all $i\in\{2,\ldots,n\}$. By Theorem 3.2, the runtime for inserting $a_{\pi(i)}$ using this pointer prediction is $O(\log\vec{\eta}(a_{\pi(i)},\mathcal{Q}^{i-1}))$. The expected total time for inserting all the elements into the priority queue is therefore at most proportional to